OpenStudy (anonymous):

the area of a rectangular window is to be 306 square centimeters. if the length exceeds width by 1 centimeter what are the demensions? how would you put this into quadratic form

hero (hero):

I would help you, but I've already goofed up twice trying to.

hero (hero):

The ironic thing is, I goofed up on stuff I'm supposed to be good at

OpenStudy (anonymous):

i mean is x+(x+1)=306^2 set up right?

OpenStudy (anonymous):

L w=306, L=w+1 w^2+w-306=0 (w-17) (w+18)=0 w=17, L=18 17*18 = 306